### v2

# Oligomorphic groups, categories of partial bijections, and ultrahomogeneous cubic spaces over finite fields

Oligomorphic groups, categories of partial bijections, and ultrahomogeneous cubic spaces over finite fields
Yury A. Neretin [^1] The work is supported by the grants of FWF (the Austrian Scientific Funds), P31591, PAT5335224.
We show that for certain class of oligomorphic groups there is a version of multiplication of double cosets in the Ismagilov–Olshanski sense. Categories of (reduced) double cosets are realized as certain categories of partial bijections. As an example, we consider the ultrahomogeneous linear space $\mathcal{Z}$ of countable dimension over a finite field equipped with a cubic form. The group of automorphisms of $\mathcal{Z}$ is an oligomorphic group, we describe its open subgroups. According Tsankov, this gives a classification of its unitary representations. The category of reduced double cosets in this case is the category of partial linear bijections of finite-dimensional cubic spaces preserving cubic forms.
0a Introduction
0.1. Categories of double cosets. In theory of unitary representations of ’large’ (’infinite-dimensional’ groups [^2] A. M. Vershik proposed the term ’large groups’ instead of ’infinite-dimensional groups’. Numerous ’infinite-dimensional groups’ are totally disconnected, so they have dimension 0 as topological spaces. This is so for groups of types $2^{\circ}$ and $4^{\circ}$-$7^{\circ}$ below. We emphasize that infinite dimensional unitary groups and (zero-dimensional) infinite symmetric groups are close relatives (at least from the point of view of representation theory).), quite often the following phenomena arise. Let $G$ be a group, and $K$ is a subgroup. It can happen that the space of double cosets $K\backslash G/K$ admits a natural associative multiplication
$K\backslash G/K\,\times\,K\backslash G/K\,\to\,K\backslash G/K,$
and the semigroup $K\backslash G/K$ acts in unitary representations of $G$.
Initially, such semigroups appeared as tools for classification of unitary representations and descriptions of spherical functions. The first examples were discovered by Rais Ismagilov [14], [15]. A wider classes of such multiplications appeared in works by Grigori Olshanski (see [35][38]). Now it is clear that this phenomenon is widespread. In particular, it arises for the following types of groups:
$1^{\circ}.$ infinite-dimensional real classical groups, see, for example, [37], [38], [25], [26];
$2^{\circ}.$ various versions of ’infinite symmetric groups’, see, for example, [35], [36], [38], [25], [27];
$3^{\circ}.$ groups of transformations of measure spaces, see, for example, [25], [32];
$4^{\circ}.$ infinite-dimensional classical groups over finite fields, see, for example, [38], [29];
$5^{\circ}.$ infinite-dimensional groups over non-Archimedean locally compact fields, see, for example, [30];
$6^{\circ}.$ groups of automorphisms and spheromorphisms of non-locally finite trees and ${\mathbb{R}}$-trees, see, for example, [34], [31].
Remarks. a) In all cases, $K$ looks as an infinite-dimensional counterpart of compact groups. In [25], Chapter VIII, there was proposed a heuristic notion of a ’heavy group’. It turns out to be that for groups $G$ with a heavy subgroup $K$ the multiplication of double cosets arises more or less automatically [^3] Examples of heavy groups are the complete symmetric group $S(\infty)$, the complete infinite hyperoctahedral group, the complete infinite-dimensional unitary group $\mathrm{U}(\infty)$, the complete orthogonal group $\mathrm{O}(\infty)$, the group of measure preserving transformations of a Lebesgue measure space. Oligomorphic groups discussed below look like ’heavy groups’.
All such groups $K$ are Roelcke precompact. This means that for any open subset $U$ in $K$ there is a finite number of elements $h_{j}$ such that the sets $Uh_{j}U$ cover $K$ (see, e.g., [40], [4]). All ’heavy groups’, that appeared in such representation-theoretical considerations are Roelcke precompact. But according [3], the group of all non-singular transformations of a Lebesgue measure space is Roelcke precompact, and it looks completely differently.
b) For groups of matrices over finite fields and local fields existing picture is relatively poor (comparing $1^{\circ}$$2^{\circ}$). But it is non-trivial and drastically differs from the real case. $\boxtimes$
It turns out that in all examined cases, we actually have families of subgroups $K_{\alpha}$ enumerated by elements $\alpha$ of some set $A$. We have a collection of operations
$K_{\alpha}\backslash K/K_{\beta}\,\times K_{\beta}\backslash K/K_{\gamma}\,\to\,K_{\alpha}\backslash K/K_{\gamma}.$
In fact, we come to a category (train of $G$), whose objects are enumerated by elements $\alpha\in A$, sets of morphisms $\beta\to\alpha$ are $K_{\alpha}\backslash K/K_{\beta}$, and products of morphisms are products of double cosets.
0.2. Oligomorphic groups. The first purpose of the present paper is to add the line:
$7^{\circ}.$ oligomorphic groups
to the list $1^{\circ}$$6^{\circ}$ above.
Denote by $S(\Omega)$ the group of all permutations of a countable space $\Omega$. An oligomorphic group (see [8]) is a closed subgroup in $S(\Omega)$, which has a finite number of orbits on any finite product $\Omega\times\dots\times\Omega$. See a collection of examples below in Subsect. 2.
This notion arose in model theory (which is a branch of mathematical logic). It is interesting to link this topic with other investigations in large groups. Some oligomorphic groups look exotic from the point of view of representation theory, but some of them are important (or at least look naturally). Todor Tsankov [40] proved that all unitary representations of oligomorphic groups are induced from finite-dimensional representations of open subgroups.
This statement allows to easily describe multiplications of reduced double cosets for oligomorphic groups (under some aditional conditions), the corresponding categories are certain categories of partial bijections (the term ’reduced double cosets’ is explained below in Subsect. 1).
The work [40] was used earlier in [29], [30]. Since oligomorphic groups look like heavy groups, the next (open) question is a search of interesting overgroups of oligomorphic groups, see a discussion in Subsect. 2.
0.3. Universal cubic spaces and groups of their automorphisms. The second purpose of the paper is to obtain a classification of unitary representations of the group of automorphisms of a universal cubic space over a finite field.
Consider a finite or countable field $𝕜\Bbbk$. We say that a cubic space is a linear space $Z$ over $𝕜\Bbbk$ equipped with a symmetric trilinear form $𝕜Z\times Z\times Z\to\Bbbk$. A linear space $𝕜\mathcal{Z}(\Bbbk)$ over $𝕜\Bbbk$ of countable dimension is universal if any finite-dimensional cubic space $X$ over $𝕜\Bbbk$ admits an embedding to $𝕜\mathcal{Z}(\Bbbk)$, and the set of such embeddings is a homogeneous space with respect to the group $𝕜\mathop{\mathrm{Isoc}}\nolimits\bigl{(}\mathcal{Z}(\Bbbk)\bigr{)}$ of automorphisms of $𝕜\mathcal{Z}(\Bbbk)$. Such space exists [^4] Universal cubic spaces (and more generally universal tensor spaces [13]) are representatives of numerous counterparts of the Urysohn universal metric spaces [41] (as the Rado graph, the Hall universal group, the universal poset, the Gurarii space, ultrametric Urysohn spaces, etc.), see, e.g., [24]. and is unique. It was a topic of numerous investigations in last years, see, for example, [5][6], [17], [13].
If a field $𝕜\Bbbk$ is finite, then the group $𝕜\mathop{\mathrm{Isoc}}\nolimits\bigl{(}\mathcal{Z}(\Bbbk)\bigr{)}$ is oligomorphic. We classify its open subgroups and unitary representations. The objects of the corresponding category are finite dimensional cubic spaces over $𝕜\Bbbk$, a morphism $V\to W$ is a form preserving invertible linear operator from a subspace $A\subset X$ to a subspace $B\subset Y$.
1. Multiplications of double cosets and trains of ’large’ groups
1.1. Multiplication of double cosets. Let $G$ be a group, $K$, $K_{1}$, $K_{2}$ be its subgroups. Denote by $K_{1}\backslash G/K_{2}$ the set of double cosets of $G$ with respect to $K_{1}$ and $K_{2}$, i.e., the quotient of $G$ by the equivalence relation
$g\sim h_{1}gh_{2},\quad\text{where $h_{1}\in K_{1}$, $h_{2}\in K_{2}$.}$
We denote classes of equivalence (points of the quotient) as $K_{1}gK_{2}$.
Let $\rho$ be a unitary representation of a separable topological group $G$ in a separable Hilbert space $\mathcal{L}$. Denote by $\mathcal{L}^{K}\subset\mathcal{L}$ the subspace of $K$-fixed vectors. By $P^{K}$ we denote the orthogonal projection to $\mathcal{L}^{K}$. It is easy to see that an operator
$\widehat{\rho}(g):=P^{K}\rho(g)\,P^{K}$
depends only on the double coset $\mathfrak{g}=KgK$ containing $g$. Indeed, let $k_{1}$, $k_{2}\in K$, $\xi$, $\eta\in\mathcal{L}$. Then matrix elements of operators $P^{K}\rho(k_{1}gk_{2})P^{K}$ and $P^{K}\rho(g)P^{K}$ coincide. Indeed,
$\bigl{\langle}P^{K}\rho(k_{1}gk_{2})P^{K}\xi,\,\eta\bigr{\rangle}_{\mathcal{L}}=\bigl{\langle}P^{K}\rho(k_{1})\rho(g)\rho(k_{2})P^{K}\xi,\,\eta\bigr{\rangle}_{\mathcal{L}}=\\ =\bigl{\langle}P^{K}\rho(k_{1})\rho(g)P^{K}\xi,\,\eta\bigr{\rangle}_{\mathcal{L}}=\bigl{\langle}\rho(k_{1})\rho(g)P^{K}\xi,\,P^{K}\eta\bigr{\rangle}_{\mathcal{L}}=\\ =\bigl{\langle}\rho(g)P^{K}\xi,\,\rho(k_{1}^{-1})P^{K}\eta\bigr{\rangle}_{\mathcal{L}}=\bigl{\langle}\rho(g)P^{K}\xi,\,P^{K}\eta\bigr{\rangle}_{\mathcal{L}}=\bigl{\langle}P^{K}\rho(g)P^{K}\xi,\,\eta\bigr{\rangle}_{\mathcal{L}}.$
Remark. Decompose $\mathcal{L}=\mathcal{L}^{K}\oplus(\mathcal{L}^{K})^{\bot}$. Then the operator $\widehat{\rho}(g)$ has a block form $\begin{pmatrix}\widetilde{\rho}(g)&0\\ 0&0\end{pmatrix}$, where $\widetilde{\rho}(g)$ is an operator in $\mathcal{L}^{K}$. $\boxtimes$
If a group $G$ is finite, then for certain non-negative rational constants $c_{\vphantom{|}\mathfrak{p}\,\mathfrak{q}}^{\vphantom{|}\mathfrak{r}}$ we have
(1.1) $\widehat{\rho}(\mathfrak{p})\,\widehat{\rho}(\mathfrak{q})=\sum_{\mathfrak{g}\in K\setminus G/K}c_{\vphantom{|}\mathfrak{p}\,\mathfrak{q}}^{\vphantom{|}\mathfrak{r}}\,\,\widehat{\rho}(\mathfrak{r})$
for any unitary representation $\rho$ of $G$.
Equivalently, let us consider the group algebra ${\mathbb{C}}[G]$ of $G$, denote $\varphi\star\psi$ the convolution in ${\mathbb{C}}[G]$. For $p\in G$ denote by $\delta_{p}\in{\mathbb{C}}[G]$ the function, which is 1 at $p$ and zero at all other points, so $\delta_{p}\star\delta_{q}=\delta_{pq}$.
Consider the subalgebra ${\mathbb{C}}[K\setminus G/K]\subset{\mathbb{C}}[G]$ consisting of $K$-biinvariant functions $\varphi$:
$\varphi(k_{1}gk_{2})=\varphi(g)\qquad\text{for any $g\in G$, $k_{1}$, $k_{2}\in K$.}$
For each double coset $\mathfrak{p}$, we consider the ’delta-function’ of $\mathfrak{p}$:
$\Delta_{\mathfrak{p}}(g)=\frac{1}{\#\mathfrak{p}}\sum_{g\in\mathfrak{p}}\delta_{g}=\begin{cases}\frac{1}{\#\mathfrak{p}},\quad&\text{if $g\in\mathfrak{p}$};\\ 0,\quad&\text{otherwise},\end{cases}$
where #$\#\mathfrak{p}$ denotes the number of elements of $\mathfrak{p}$. Such functions form a basis in ${\mathbb{C}}[K\setminus G/K]$. The function $\Delta_{K}:=\Delta_{KeK}$ is an idempotent in ${\mathbb{C}}[G]$ and the unit in ${\mathbb{C}}[K\setminus G/K]$:
$\Delta_{K}\star\Delta_{K}=\Delta_{K},\quad\Delta_{K}\star\delta_{p}\star\Delta_{K}=\Delta_{KpK},\quad\Delta_{K}\star\Delta_{\mathfrak{p}}=\Delta_{\mathfrak{p}}=\Delta_{\mathfrak{p}}\star\Delta_{K}.$
For any unitary representation $\rho$ of $G$ we have
$\rho\bigl{(}\Delta_{K}\bigr{)}=P^{K},$
and therefore,
$\rho\bigl{(}\Delta_{KpK}\bigr{)}=\rho\bigl{(}\Delta_{K}\star\delta_{p}\star\Delta_{K}\bigr{)}=\rho\bigl{(}\Delta_{K}\bigr{)}\,\rho\bigl{(}\delta_{p})\,\rho\bigl{(}\Delta_{K}\bigr{)}=P^{K}\,\rho\bigl{(}p)\,P^{K}=\widehat{\rho}(\mathfrak{p}).$
So, the coefficients in (1.1) are simply the structure constants of the convolution algebra ${\mathbb{C}}[K\backslash G/K]$,
(1.2) $\Delta_{\mathfrak{p}}\star\Delta_{\mathfrak{q}}=\sum_{\mathfrak{g}\in K\setminus G/K}c_{\vphantom{|}\mathfrak{p}\,\mathfrak{q}}^{\vphantom{|}\mathfrak{r}}\,\,\Delta_{\mathfrak{r}}.$
If $G$ is locally compact and $K$ is a compact subgroup, then there is a system of probabilistic measures $\mu_{\vphantom{|^{|}}\mathfrak{p}\,\mathfrak{q}}$ on $K\setminus G/K$ such that
[…]
for all unitary representations $\rho$.
Equivalently, we can consider convolution algebras of $K$-biinvariant functions (or $K$-biinvariant measures) on $G$.
Such ’Hecke–Iwahori algebras’ of double cosets were widely used in classical representation theory. But except few examples (including Hecke algebras, affine Hecke algebras [^5] Both objects and names for them were introduced by Nagayoshi Iwahori., and closely related Yokonuma algebras) algebras of double cosets are heavy objects.
It turns out that for a ’large group’ $G$ and its subgroup $K$ quite often there exists a natural associative multiplication $(\mathfrak{p},\mathfrak{q})\mapsto\mathfrak{p}\circ\mathfrak{q}$ on the spae $K\backslash G/K$ such that
(1.4) $\widehat{\rho}(\mathfrak{p})\,\widehat{\rho}(\mathfrak{q})=\widehat{\rho}(\mathfrak{p}\circ\mathfrak{q})$
for all unitary representations $\rho$ of the group $G$.
Remark. Olshanski [35], [37], [38] proposed the following explanation of this phenomenon. Let $G_{n}$ be finite-dimensional (or finite) groups, let $G_{\infty}$ be their ’limit’ (as symmetric groups $S_{n}$ and $S_{\infty}$). Then, in formulas of type (1.2), we get a concentration of convolutions near one double coset (but this explanation does not always works; moreover, generally, this does not formally imply (1.4)). $\boxtimes$
It is important that such multiplications $\mathfrak{p}\circ\mathfrak{q}$ admit transparent descriptions (in a contrary with formulas of the type (1.1)–(1.3)).
1.2. Reduced double cosets. Generally, functions $\mathfrak{g}\mapsto\widehat{\rho}(\mathfrak{g})$, where $\rho$ ranges in the space of all unitary representations, do not separate elements $\mathfrak{g}$ of $K\backslash G/K$. Definitely, this is so if the topological quotient space $K\backslash G/K$ is non-Hausdorff (orbit spaces of non-compact groups are non-Hausdorff very often). Olshanski ([38], Sect. 4) found an example, when this happens for finite Hausdorff spaces $K\backslash G/K$.
Consider a separable topological group $G$, its subgroup $K$, and a family of subgroups $K_{\alpha}\subset K$ enumerated by a set $\mathcal{A}$ (and this family includes $K$). For any unitary representation $\rho$ of $G$ in a Hilbert apace $\mathcal{L}$, we denote by $\mathcal{L}_{\alpha}$ the subspace of $K_{\alpha}$-fixed vectors, by $P_{\alpha}$ the orthogonal projection to $\mathcal{L}_{\alpha}$. Consider operators
$\widetilde{\rho}_{\alpha,\beta}(g):=P_{\alpha}\,\rho(g)\Bigr{|}_{\mathcal{L}_{\beta}}:\,\mathcal{L}_{\beta}\to\mathcal{L}_{\alpha}.$
We define reduced double cosets $[\![K_{\alpha}\backslash G/K_{\beta}]\!]$ as quotients of $G$ by the following equivalence relation:
$g_{1}\sim g_{2}\quad\text{if $\widetilde{\rho}(g_{1})=\widetilde{\rho}(g_{2})$ for all unitary representations $\rho$.}$
We say that $G$ is a $(G,K)$-pair if for each $\alpha$, $\beta$, $\gamma\in\mathcal{A}$ for any $\mathfrak{p}\in[\![K_{\alpha}\backslash G/K_{\beta}]\!]$, $\mathfrak{q}\in[\![K_{\beta}\backslash G/K_{\gamma}]\!]$ there exists an element $\mathfrak{p}\circ\mathfrak{q}\in[\![K_{\alpha}\backslash G/K_{\gamma}]\!]$ such that the following property of multiplicativity holds:
(1.5) $\widetilde{\rho}_{\alpha,\beta}(\mathfrak{p})\,\widetilde{\rho}_{\beta,\gamma}(\mathfrak{q})=\widetilde{\rho}_{\alpha,\gamma}(\mathfrak{p}\circ\mathfrak{q})$
for all unitary representations $\rho$ of $G$.
Since a product of linear operators is an associative operation, the product $\mathfrak{p}\circ\mathfrak{q}$ also is associative. So we come to a category $\operatorname{Tra}\bigl{(}G,\{K_{\alpha}\}\bigr{)}$ (the train of $G$). Its objects are enumerated by elements of $\mathcal{A}$. Morphisms $\beta\to\alpha$ are elements of $[\![K_{\alpha}\backslash G/K_{\beta}]\!]$. A product of morphisms is the $\circ$-product.
This category is equipped with the involution $\mathfrak{p}\mapsto\mathfrak{p}^{*}$ acting from $[\![K_{\alpha}\backslash G/K_{\beta}]\!]$ to $[\![K_{\beta}\backslash G/K_{\alpha}]\!]$ for each pair $\alpha$, $\beta$, it is determined by the map $g\mapsto g^{-1}$. We have
$(\mathfrak{p}\circ\mathfrak{q})^{*}=\mathfrak{q}^{*}\circ\mathfrak{p}^{*},\qquad\mathfrak{p}^{**}=\mathfrak{p}.$
By the definition, any unitary representation $\rho$ of the group $G$ generates a representation $\widetilde{\rho}$ of the category $\operatorname{Tra}(G,\{K_{\alpha}\})$ (for a formal definition of representations of categories, see the next subsection).
1.3. Representations of categories. Recall basic definitions and notation related to representations of categories. For details, discussion, and further definitions, see [25], Sect. II.8.
Let $\mathcal{K}$ be a category. Denote by $\operatorname{Ob}(\mathcal{K})$ the set of objects of $\mathcal{K}$, by $\operatorname{Mor}(V,W)=\operatorname{Mor}_{\vphantom{|^{|}}\mathcal{K}}(V,W)$ the set all morphisms $V\to W$, by $\operatorname{End}(V)$ the semigroup of all endomorphisms of an object $V$. We assume that these semigroups have units $1_{V}$ and these units are units in the category $\mathcal{K}$, i.e., for any $\mathfrak{p}\in\operatorname{Mor}(V,W)$ we have
$\mathfrak{p}\cdot 1_{V}=1_{W}\cdot\mathfrak{p}=\mathfrak{p}.$
Denote by $\operatorname{Aut}(V)$ the group of automorphisms of an object $V$. An involution in $\mathcal{K}$ is a collection of maps $\mathfrak{p}\mapsto\mathfrak{p}^{*}$ acting from $\operatorname{Mor}(V,W)$ to $\operatorname{Mor}(W,V)$ for all $V$, $W$ such that $(\mathfrak{p}\mathfrak{q})^{*}=\mathfrak{q}^{*}\mathfrak{p}^{*}$, $\mathfrak{p}^{**}=\mathfrak{p}$.
A *-representation $(Q,\varkappa)$ of a category $\mathcal{K}$ is a functor from $\mathcal{K}$ to the category of Hilbert spaces and bounded operators. Namely, for each $V\in\operatorname{Ob}(\mathcal{K})$ we assign a Hilbert space $Q(V)$ and for each $\mathfrak{p}\in\operatorname{Mor}(V,W)$ we assign a bounded operator
$\varkappa_{\vphantom{|^{|}}W,V}(\mathfrak{p}):Q(V)\to Q(W)$
such that for any $V$, $W$, $Y\in\operatorname{Ob}(\mathcal{K})$ and for any $\mathfrak{p}\in\operatorname{Mor}(V,W)$, $\mathfrak{q}\in\operatorname{Mor}(W,Y)$ we have
[…]
We also require
[…]
1.4. Example: the infinite symmetric group $S(\infty)$ and the category of partial bijections. See Olshanski [35], see also [25], Sect. VIII.1-2. Let $\Omega$ be a countable or finite set, $S(\Omega)$ be the group of all permutations of $\Omega$. Denote $S_{\infty}:=S({\mathbb{N}})$. It is known, that this group (infinite symmetric group) has a unique separable topology (see [18]). Namely, $g_{j}$ converges to $g$ if for each $\omega\in\Omega$ we have $g_{j}\omega=g\omega$ for sufficiently large $j$.
Let $G=K=S_{\infty}$. Denote by $K_{\alpha}\subset S_{\infty}$ the subgroup consisting of elements fixing $1$, …, $\alpha\in{\mathbb{N}}$, we assume $K_{0}:=S_{\infty}$. By $K^{\circ}_{\alpha}$ we denote the subgroup in $K$ consisting of elements sending the set $\{1,\dots,\alpha\}$ to itself. The quotient group $K^{\circ}_{\alpha}/K_{\alpha}$ is the finite symmetric group $S_{\alpha}$.
Arthur Lieberman [20] obtained the classification of all unitary representations of $S_{\infty}$. Any irreducible representation of this group is induced from an irreducible representation of some subgroup $K^{\circ}_{\alpha}$ trivial on $K_{\alpha}$.
The train of $S_{\infty}$. Let $g\in S_{\infty}$. Denote by $(g)_{\alpha\beta}$ the partial bijection $\{1,\dots,\beta\}\to\{1,\dots,\alpha\}$, which sends $i\leqslant\beta$ to $j\leqslant\alpha$ if and only if $g$ sends $i$ to $j$. It is easy to show that the map $g\mapsto(g)_{\alpha,\beta}$ determines a bijection between double cosets $K_{\alpha}\backslash S_{\infty}/K_{\beta}$ and partial bijections $\{1,\dots,\beta\}\to\{1,\dots,\alpha\}$.
Recall the definition of partial bijections (see, e.g., [25], Sect. VIII.1). Let $X$, $Y$ be sets. A partial bijection $\mu:X\to Y$ is a bijection from a subset $A\subset X$ to a subset $B\subset Y$. We say that $A$ is the domain $\operatorname{dom}\mu$ of $\mu$, $B$ is the image $\operatorname{im}\mu$. A product of partial bijections $\mu:X\to Y$, $\nu:Y\to Z$ is defined in the following way: $\operatorname{dom}\nu\mu:=\mu^{-1}(\operatorname{dom}\nu\,\cap\,\operatorname{im}\mu)$ and $\nu\mu(x)=\nu(\mu(x))$ for $x\in\operatorname{dom}\nu\mu$. We also define the pseudoinverse partial bijection, if $\mu:X\to Y$ is as above, then its pseudoinverse $\mu^{*}:X\to Y$ is the inverse bijection $B\to A$.
Denote by ${\operatorname{PB}}$ the category of partial bijections of finite sets. According [35] (see also exposition in [25], Sect. VIII.1-2), the following theorem holds:
Theorem 1.1.
a) The category ${\operatorname{PB}}$ is the train $\operatorname{Tra}\bigl{(}S_{\infty},\{K_{\alpha}\}\bigr{)}$.
b) The functor $\rho\mapsto\widetilde{\rho}_{\alpha,\beta}$ is a one-to-one correspondence between unitary representations of $S_{\infty}$ and *-representations of the category ${\operatorname{PB}}$.
Classification of *-representations of the category ${\operatorname{PB}}$ is a simple problem, see [25], Sect. VIII.1-2 (also, see below, Subsect. 2). In fact, Theorem 1.1 is equivalent to Lieberman’s description [20] of unitary representations of $S_{\infty}$.
2. Oligomorphic groups and their trains
Here we construct categories of reduced double cosets for oligomorphic groups. This is a simple corollary of description of unitary representations of oligomorphic groups obtained by Tsankov [40]. These categories embed in the semigroups of partial bijections examined by Ben Yaacov, Ibarlucía, Tsankov [4].
2.1. Oligomorphic groups. A closed subgroup $G$ in an infinite symmetric group $S(\Omega)$ is called oligomorphic if it has a finite number of orbits on each multiple product $\Omega^{n}:=\Omega\times\dots\times\Omega$ ($n$ times), see Cameron [8]. See a collection of examples below in Subsect. 2.
The stabilizer of any finite subset in $\Omega$ is oligomorphic.
If $H\subset G$ is an open subgroup, then $G$ is closed in the symmetric group $S(G/H)$ and has a finite number of orbits on each set $(G/H)^{n}$.
Tsankov [40] obtained the following description of unitary representations of such groups:
Theorem 2.1.
a) Each representation of $G$ is a direct sum of irreducible representations [^6] For this reason, the group $G$ has type $I$ (see, e.g., [23], Chaper 1)..
b) For any irreducible representation $\rho$ of $G$ there is an open subgroup $H^{\circ}$ and a subgroup $H\subset H^{\circ}$ of finite index such that $\rho$ is inducedl [^7] For induced representations of totally disconnected groups, see Mackey [21], [23]. Also, see below Subsect. 2. from a representation of $H^{\circ}$ trivial on $H$.
Irreducible representations induced from subgroups $H$ and $H^{\prime}$ can be equivalent only for trivial reasons: if $H$ and $H^{\prime}$ are conjugate and their representations $\sigma$, $\sigma^{\prime}$ are conjugate.
2.2. Perfect oligomorphic groups. Fix an oligomorphic subgroup $K\subset S(\Omega)$. For a finite subset $A\subset\Omega$ we denote by $K(A)\subset S(\Omega)$ the subgroup consisting of all $g\in S(\Omega)$ fixing $A$ pointwise. We say that a finite subset $A\subset\Omega$ is perfect [^8] In the terminology of model theory, ’a perfect set’ is ’an algebraically closed set’ (origin of this words is clear from Examples $3^{\lozenge}$, $4^{\lozenge}$ of Subsect. 2 or Proposition 3.4), see [11], [40]. ’Perfect oligomorphic groups’ are ’groups of automorphisms of first order structures admitting a weak elimination of singularities’, See also [11], Lemma 1.3. if all orbits of $K(A)$ on $\Omega\setminus A$ are infinite. For a perfect $A$ denote by $K^{\circ}(A)\subset S(\Omega)$ the subgroup sending $A$ to itself. Denote
$\Gamma(A):=K^{\circ}(A)/K(A),$
it is a finite group acting on $A$.
Lemma 2.2.
a) Let $A\subset\Omega$ be perfect, $g\in K$. Then $gA$ is perfect.
b) If $A$, $B$ are perfect, then $A\cap B$ is perfect.
c) For any finite $R\subset\Omega$ there is a minimal perfect set $\overline{\overline{R}}$ containing $R$.
Proof. Statements a), b) are obvious.
Since $K$ is oligomorphic, the group $K(R)$ has a finite number of orbits on $\Omega$. In particular, the number of finite orbits is finite. Let $\overline{\overline{R}}$ be the union of all finite orbits of $K(R)$. The group $K(\overline{\overline{R}})$ has finite index in $K(R)$ and elements of $\Omega\setminus\overline{\overline{R}}$, which have finite orbits with respect to $K(\overline{\overline{R}})$, have finite orbits with respect to $K(R)$. So such elements are absent. $\square$
Notice that the set of classes of $K$-equivalent perfect sets has a natural partial order:
$B\preceq C$
if there is $g\in K$ such that $gB\subset C$.
We say that a closed subgroup $G$ in $S(\Omega)$ is perfect oligomorphic if it is oligomorphic and for each open subgroup $H\subset G$ there is a perfect finite subset $A$ such that $K(A)\subset H\subset K^{\circ}(A)$.
Remark. This property depends on an embedding of $G$ in the infinite symmetric group. For instance, consider the set $\Omega^{[k]}$ of all $k$-element subsets in $\Omega$ ($k>1$) and the action of $S(\Omega)$ on $\Omega^{[k]}$. Clearly, the subgroup $S(\Omega)$ in $S\bigl{(}\Omega^{[k]}\bigr{)}$ is not perfect oligomorphic. $\boxtimes$
Fix a perfect oligomorphic group $K\subset S(\Omega)$. Let $A$, $B$ be perfect subsets in $\Omega$. We say that a partial bijection $\theta:A\to B$ is a restriction $g\bigr{|}_{A\to B}$ of $g\in K$ if $\operatorname{dom}\theta=A\cap g^{-1}B$ and the bijection $\operatorname{dom}\theta\to\operatorname{im}\theta$ is induced by $g$.
Lemma 2.3.
Let $K\subset S(\Omega)$ be perfect oligomorphic. Let $A$, $B$, $C\subset\Omega$ be perfect sets. Let partial bijections $\mu:A\to B$ , $\nu:B\to C$ have the form
$\mu=p\Bigr{|}_{A\to B},\quad\nu=q\Bigr{|}_{B\to C},\quad\text{where $q$, $p\in K$.}$
Then there is $r\in K$ such that $\nu\mu=r\Bigr{|}_{A\to C}$.
We need the following P. Neumann Lemma (see [33], [8], (2.16)):
Lemma 2.4.
Let a group $G$ act on a countable set $\Xi$ without finite orbits. Let $M$, $N$ be finite subsets in $\Xi$. Then there is $g\in G$ such that $gM\cap N=\varnothing$.
Proof of Lemma 2.3. We set
$\Xi:=\Omega\setminus B,\quad G:=K(B),\quad M:=p(A)\setminus B,\quad N:=q^{-1}B\setminus C$
and choose $g$ as in the Neumann Lemma. We set $r=qgp$. $\square$
2.3. Trains of perfect oligomorphic groups. Consider the following category $\operatorname{Tra}(K)$ (the train of a perfect oligomorphic group $K\subset S(A)$). Its objects are perfect finite subsets in $\Omega$. A morphism $\theta:A\to B$ is a partial bijection that can be obtained by a restriction of $g\in K$ to $A$. A product of morphisms is the product of partial bijections (by Lemma 2.3 such product is a morphism). The involution in $\operatorname{Tra}(K)$ is the involution in the category of partial bijections.
Remark. So $\operatorname{Tra}(K)$ is realized by partial bijection and the set of morphisms is closed with respect to the involution. Therefore, it is an ’inverse category’ in the sense of [16], i.e., a categorical counterpart of inverse semigroups. $\boxtimes$
Theorem 2.5.
Let $K\subset S(\Omega)$ be a perfect oligomorphic group, let $A$ range in the space of perfect subsets in $\Omega$.
a) The category $\operatorname{Tra}(K)$ is the train $\operatorname{Tra}(K,\{K(A)\})$ of the group $K$.
b) The map $\rho\mapsto\widetilde{\rho}$ is a bijection between unitary representations of $K$ and *-representations of the category $\operatorname{Tra}(K)$.
Remark. We also can reduce $\operatorname{Tra}(K)$ choosing one representative from each $K$-equivalence class of perfect sets. In this way, we pass to an equivalent category, which is ’smaller’ and more transparent. $\boxtimes$
The proof of this statement given below in Subsect. 22 is a simple corollary of Tsankov’s Theorem 2.1, formal manipulations with ’ordered categories’ in the sense of [25], and Mackey’s technique of induced representations [21][23], [2].
2.4. Description of representations of $\operatorname{Tra}(K)$ corresponding to representations of $K$. Here we verify the statement a) of Theorem 2.5.
According Theorem 2.1, any irreducible representation of a perfect oligomorphic group $K$ is induced by an irreducible representation of some subgroup $K^{\circ}(A)$ trivial on $K(A)$.
Let us reformulate a definition of induced representations [21] in a convenient for us form. Let $A\subset\Omega$ be a perfect set. Denote by $\mathcal{E}(A)$ the set of all maps $\iota:A\to\Omega$ obtained by restrictions of elements $g\in K$. Clearly, $\mathcal{E}(A)$ is a $K$-homogeneous space $K/K(A)$, the group $K$ acts on $\mathcal{E}(A)$ by left multiplications $\iota\mapsto g^{-1}\iota$. We also have an action of the finite group $\Gamma(A)=K^{\circ}(A)/K(A)$ on $\mathcal{E}(A)$ by right multiplications $\iota\mapsto\iota\gamma$. These two actions commute.
Consider a unitary irreducible representation $\sigma$ of $\Gamma(A)$ in a Euclidean space $V$. Consider the space $\ell^{2}(\mathcal{E}(A),V)$ of $V$-valued functions on $\mathcal{E}(A)$ equipped with the $\ell^{2}$-inner product
$\langle F_{1},F_{2}\rangle_{\ell^{2}}=\sum_{\iota\in\mathcal{E}\,(A)}\langle F_{1}(\iota),F_{2}(\iota)\rangle_{V}.$
The group $K$ acts in $\ell^{2}(\mathcal{E}(A),V)$ by shifts
(2.1) $\pi(g)F(\iota)=F(g^{-1}\iota).$
We consider the subspace $H=H(A,\sigma)\subset\ell^{2}(\mathcal{E}(A),V)$ consisting of functions satisfying
(2.2) $F(\iota\gamma)=\sigma(\gamma^{-1})F(\iota),$
where $\gamma\in\Gamma(K)$. Restricting operators (2.1) to $H(A,\sigma)$, we get a unitary representation
(2.3) $\tau=\operatorname{Ind}_{K^{\circ}(A)}^{K}(\sigma)$
of $K$ in $H(A,\sigma)$ (the representation of the group $K$ induced by the representation $\sigma$ of the subgroup $K^{\circ}(A)$).
Remark. So, the representation $\tau$ is a subrepresentation in the quasiregular representation of $K$ in $\ell^{2}\bigl{(}\mathcal{E}(A)\bigr{)}=\ell^{2}\bigl{(}K/K(A)\bigr{)}$. $\boxtimes$
Let us examine the operators corresponding to double cosets. For a perfect set $B$, denote by $\mathcal{E}(A,B)$ the set of all embeddings $\iota:A\to B$ such that $\iota=g\bigr{|}_{A}$ for some $g\in K$, so $\mathcal{E}(A,B)\subset\mathcal{E}(A)$. Denote by $H(A,\sigma)_{B}\subset H$ the subspace of $K(B)$-invariant functions, by $P(B)$ the projection to $H(A,\sigma)_{B}\subset H(A,\sigma)$.
Lemma 2.6.
The subspace $H(A,\sigma)_{B}$ consists of functions $F\in H(A,\sigma)$ supported by $\mathcal{E}(A,B)$. The projection $P(B)$ is the operator of restriction of functions from $\mathcal{E}(A)$ to the subset $\mathcal{E}(A,B)$.
Proof. A $K(B)$-invariant function $F\in\ell^{2}$ must be constant on $K(B)$-orbits. If $\iota(A)\not\subset B$, then $K(B)$-orbit of $\iota$ is infinite. Since $F\in\ell^{2}$, it is zero on such orbits.
If $\iota(A)\subset B$, then $\iota$ is fixed by the subgroup $K(B)$, and this does not impose additional conditions for $F$. $\square$
Fix perfect sets $B$, $C$. For any $g\in K$ we assign the partial bijection $g\Bigr{|}_{C\to B}$ as above. Clearly, the map $g\mapsto g\Bigr{|}_{C\to B}$ is constant on double cosets $K(B)\backslash K/K(C)$. So, we have a surjective (generally, non-injective) map
$K(B)\backslash K/K(C)\to\operatorname{Mor}(C,B).$
Now we define a representation $(T,\tau)$ of the category $\operatorname{Tra}(K)$. A space $T(B)$ is the space of $V$-valued functions on $\mathcal{E}(A,B)$ satisfying (2.2). An operator
$\tau_{B,C}(g)=P(B)\,\tau(g)\Bigr{|}_{T(C)}:T(C)\to T(B)$
is given by the formula:
$\tau_{B,C}(g)F(\iota)=\begin{cases}F(g^{-1}\iota),\quad&\text{if $i(A)\subset g^{-1}C\cap B$};\\ 0,\quad&\text{otherwise}.\end{cases}$
Clearly, we get a representation of the category $\operatorname{Tra}(K)$ and this establishes the statement a) of Theorem 2.5. We separately formulate the following obvious statement.
Lemma 2.7.
For the representation $(T,\tau)$ of $\operatorname{Tra}(K)$, a space $T(B)$ is nonzero if and only if $B\succeq A$. The representation of $\operatorname{End}(A)$ in $T(A)$ is zero on $\operatorname{End}(A)\setminus\operatorname{Aut}(A)$ and coincides with $\sigma$ on $\operatorname{Aut}(A)$.
2.5. Representations of the category $\operatorname{Tra}(K)$. It remains to verify that each irreducible *-representation of the category $\operatorname{Tra}(K)$ arises from an irreducible representation of $K$.
Lemma 2.8.
a) Let $(Q,\varkappa)$ be a *-representation of $\operatorname{Tra}(K)$. Let $D$ be a $\preceq$-minimal perfect set such that $Q(D)\neq 0$. Then $\varkappa_{\vphantom{|^{|}}D,D}(\varphi)=0$ for any morphism $\varphi\in\operatorname{End}(D)\setminus\operatorname{Aut}(D)$.
b) Additionally, let $(Q,\varkappa)$ be irreducible. Then the $\preceq$-minimal perfect set such that $Q(D)\neq 0$ is unique (upto $K$-equivalence).
We call this $\preceq$-minimal perfect set $D$ the conductor of an irreducible representation.
Proof. Let $\varphi\in\operatorname{End}(D)\setminus\operatorname{Aut}(D)$, $\varkappa_{D,D}(\varphi)\neq 0$. Then the partial bijection $\varphi^{*}\varphi$ is an identical map $\operatorname{dom}\varphi\to\operatorname{dom}\varphi$, and
[…]
Indeed, if $A:L_{1}\to L_{2}$ is a non-zero operator between Hilbert spaces, then $A^{*}A$ also is non-zero.
By Lemma 2.2.a)-b), $\operatorname{dom}\varphi$ is a perfect set, denote it by $C$. Let $\mu$ be the identical embedding $C\to D$, so $\mu\mu^{*}=\varphi^{*}\varphi$. Therefore, $\varkappa_{\vphantom{|^{|}}D,C}(\mu)\neq 0$. Hence,
[…]
But $\mu^{*}\mu$ is the unit in $\operatorname{End}(C)$, and therefore $Q(C)\neq 0$. We come to a contradiction.
b) Let $D$, $D^{\prime}$ be two non-equivalent $\preceq$-minimal perfect sets. Let they be non-equivalent. Let $\psi\in\operatorname{Mor}(P,P^{\prime})$. If $\operatorname{dom}\psi\neq P$, then $\psi^{*}\psi\in\operatorname{End}(P)\setminus\operatorname{Aut}(P)$. So, $\varkappa(\psi^{*}\psi)\neq 0$. Hence $\varkappa(\psi)=0$.
If $\operatorname{im}\psi\neq P^{\prime}$, then $\psi\psi^{*}\in\operatorname{End}(P^{\prime})\setminus\operatorname{Aut}(P^{\prime})$. Again, $\varkappa(\psi)=0$.
So, $\varkappa(\psi)=0$ for all $\psi\in\operatorname{Mor}(P,P^{\prime})$. This contradicts to irreducibility. $\square$
Now Theorem 2.5 follows from straightforward arguments, which are presented below.
Let $(Q,\varkappa)$ be a representation of a category $\mathcal{K}$. Let $C\in\operatorname{Ob}(\mathcal{K})$. Let $\xi_{\alpha}$ be a family of vectors in $Q(C)$. For any $D\in\operatorname{Ob}(\mathcal{K})$ we consider the closed linear span $N(D)$ of all vectors $\varkappa_{\vphantom{|^{|}}D,C}(\mu)\xi_{\alpha}$, where $\mu$ ranges in $\operatorname{Mor}(C,D)$. Clearly, the family of subspaces $N(D)$ is a subrepresentation (we call it the cyclic span of $\xi_{\alpha}$). We say that a family $\xi_{\alpha}$ is cyclic if $N(D)=Q(D)$ for all $D$.
Lemma 2.9.
A *-representation $(Q,\varkappa)$ of $\operatorname{Tra}(K)$ with a cyclic space $Q(C)$ is uniquely determined by the representation of $\operatorname{End}(C)$ in $Q(C)$.
Proof. This is a version of the ’GNS-construction’. Let $v_{1}$, $v_{2}\in Q(C)$, $r_{1}$, $r_{2}\in\operatorname{Mor}(C,D)$. Then
[…]
and we see that such inner products are uniquely determined by the representation of $\operatorname{End}(C)$. Let $(Q^{\prime},\varkappa^{\prime})$ be another representation of the category $\operatorname{Tra}(K)$, which coincides with $(Q,\varkappa)$ on $\operatorname{End}(C)$. Then for each $D$ we can identify $Q(D)$ with $Q^{\prime}(D)$ assuming that any vector $\varkappa^{\prime}_{\vphantom{|^{|}}D,C}(r)w$ corresponds to $\varkappa_{\vphantom{|^{|}}D,C}(r)w$.
[…]
Lemma 2.10.
Any *-representation $(R,\rho)$ of $\operatorname{Tra}(K)$ is a direct sum of irreducible representations.
Proof. Consider a sequence $A_{1}$, $A_{2}$, $\dots$ including all perfect sets in $\Omega$ defined up to the $K$-equivalence. Since $K$ is oligomorphic, it has only finite number of orbits on the space of finite subsets of a fixed cardinality. So, we can assume that numbers of elements of $A_{j}$ are (non-strictly) increasing. We take the cyclic span $(Q^{(1)},\varkappa^{(1)})$ of $R(A_{1})$ and the orthogonal complement $(R^{(1)},\rho^{(1)})$ of the cyclic span. By construction, we have $R^{(1)}(A_{1})=0$. Next, we take the cyclic span $(Q^{(2)},\varkappa^{(2)})$ of $R^{(1)}(A_{1})$ and orthogonal complement $(R^{(2)},\rho^{(2)})$ of $(Q^{(2)},\varkappa^{(2)})$. Now we have $R^{(2)}(A_{1})=0$, $R^{(2)}(A_{2})=0$. Etc.
In this way, we get a decomposition $(R,\rho)=\oplus_{j}(Q^{(j)},\varkappa^{(j)})$. In each summand $(Q^{(j)},\varkappa^{(j)})$, a perfect set $A_{j}$ is a minimal set for which $Q^{(j)}(\cdot)\neq 0$. By Lemma 2.8, the action of $\operatorname{End}(A_{j})$ in $Q^{(j)}(A_{j})$ is reduced to the action of $\operatorname{Aut}(A_{j})$. We decompose each representation of a finite group $\operatorname{Aut}(A_{j})$ into a direct sum, take the cyclic span of each summand with respect to $\operatorname{Tra}(K)$, and come to a decomposition of each $(Q^{(j)},\varkappa^{(j)})$ into a direct sum of irreducible representations. $\square$
According Lemmas 2.8 and 2.9, an irreducible *-representations $(R,\rho)$ of the category $\operatorname{Tra}(K)$ is uniquely determined by its conductor $A$ and an irreducible representation $\sigma$ of a finite group $\operatorname{Aut}(A)$.
According Lemma 2.7, these representations coincide with representations originated from representations of the group $K$. This completes the proof of Theorem 2.5.
2.6. Weak closures. Let $\rho$ be a unitary representation of a group $G$. Consider the image $\rho(G)$ in the unitary group and its closure $\overline{\rho(G)}$ in the space all operators with respect to the weak operator topology. It is well-known (and can be easily verified, see, for example, [38], [25], Sect. I.4) that $\overline{\rho(G)}$ is a compact metrizable semigroup with a separately continuous multiplication. See a discussion of such closures for ’infinite-dimensional’ (’large’) groups in [38], [25], for oligomorphic groups in [4].
Proposition 2.11.
Let $K\subset S(\Omega)$ be a perfect oligomorphic group. Let $B\subset\Omega$ be a perfect set. Then there is a sequence $g_{j}\in K(B)$ such that for any unitary representation $\rho$ of $K$ the orthogonal projection $P(B)$ to the subspace of $K(B)$-fixed vectors is a weak operator limit of the sequence $\rho(g_{j})$.
Remark. As far as the author knows, a similar statement holds in all cases, when multiplication of double cosets exists. $\boxtimes$
Proof. It is sufficient to verify the statement for irreducible representations. Consider an exhausting family of subsets $B\subset R_{1}\subset R_{2}\subset\dots\subset\Omega\setminus B$. By Lemma 2.4, for each $j$ there exists $g_{j}\in K(B)$ such that
$g_{j}(R_{j}\setminus B)\cap(R_{j}\setminus B)=\varnothing.$
Consider the quasiregular representation $\pi$ of $K$ in $\ell^{2}(\mathcal{E}(A,\Omega))$, see (2.1). For $\iota\in\mathcal{E}(A,\Omega)$ denote by $e_{\iota}$ the $\delta$-function supported by $\iota$. If $\iota\in\mathcal{E}(A,B)$, then $\iota$ is fixed by all $g_{j}$. Let $\iota(A)\not\subset B$. Fix $\nu\in\mathcal{E}(A)$. There is $N$ such that $(\iota(A)\setminus B)$, $(\nu(A)\setminus B)\subset R_{N}$. For large $j\geqslant N$ we have
$\bigl{(}g_{j}(\iota(A)\setminus B)\bigr{)}\cap\bigl{(}\nu(A)\setminus B\bigr{)}=\varnothing.$
Therefore, $g_{j}\iota\neq\nu$, and
$\langle\pi(g_{j})e_{\iota},e_{\nu}\rangle_{\ell^{2}}=\langle e_{g_{j}\iota},e_{\nu}\rangle_{\ell^{2}}=\delta_{g_{j}\iota,\nu}=0.$
Therefore, the weak limit of the sequence $\pi(g_{j})$ is the projection to the space of $K(B)$-fixed vectors.
Any induced representation (2.3) is a subrepresentation of $\pi$, and we get the desired statement. $\square$
Corollary 2.12.
For any unitary representation $\rho$ of a perfect oligomorphic group $K\subset\Omega$ for each perfect sets $A$, $B$ operators
$\widetilde{\rho}_{A,B}(g)=P(A)\,\rho(g)P(B)$
are contained in the weak closure of $\rho(K)$.
So, the category $\operatorname{Tra}(K)$ has a canonical embedding in the universal universal semigroup compactification of $K$, which was examined by Ben Yaacov, Ibarlucía, Tsankov [4].
2.7. Examples. We discuss an initial collection of examples of oligomorphic groups (for $1^{\lozenge}$-$4^{\lozenge}$, see [40], Sect. 5, and the corresponding categories.
$0^{\lozenge}.$ Infinite symmetric group, see our Subsect. 1.
$1^{\lozenge}.$ Consider the group $G=\operatorname{Aut}({\mathbb{Q}},<)$ of all order preserving bijections of the set ${\mathbb{Q}}$ of rational numbers. Perfect sets are arbitrary finite subsets $A\subset{\mathbb{Q}}$. We come to the category $\operatorname{Tra}\bigl{(}\operatorname{Aut}({\mathbb{Q}},<)\bigr{)}$. It objects are finite ordered sets. Morphisms $A\to B$ are order preserving partial bijections. Groups of automorphisms are trivial.
So, irreducible unitary representations are enumerated by the number #$\#A$ of elements of a conductor $A$. In other words, any irreducible representations of this group is realized in the spaces $\ell^{2}$ on the set of all $k$-element subsets in ${\mathbb{Q}}$ for some $k$.
$2^{\lozenge}.$ Let $\mathcal{R}$ be the Rado graph, see, e.g., [8]. Recall that it is a graph with a countable number of vertices satisfying the condition (the ’Urysohn extension property’): let $A\subset B$ be finite graphs, then any embedding $A\to\mathcal{R}$ extends to an embedding $B\to\mathcal{R}$.
Consider a countable set, call its elements ’vertices’, for each pair of vertices draw an edge connecting them with probability $1/2$. Then we get the Rado graph with probability 1, see [10], [9].
The group $\operatorname{Aut}(\mathcal{R})$ of automorphisms of $\mathcal{R}$ acting on the set of vertices is perfect oligomorphic, arbitrary finite subset of the set of vertices is perfect. Objects of the train $\operatorname{Tra}(\operatorname{Aut}(\mathcal{R}))$ are finite graphs, a morphism $X\to Y$ is an isomorphism of an induced subgraph [^9] To avoid a misleading we use the term ’induced subgraph’: we consider some set of vertices of $X$ and all edges of $X$ connecting these vertices. $A\subset X$ to an induced subgraph $B\subset Y$.
Irreducible representations are enumerated by pairs
(a finite graph Δ defined up to isomorphism,an irreducible representation of the group Aut(Δ) defined up to $\Bigl{(}\text{a finite graph $\Delta$ defined up to isomorphism,}\\ \text{an irreducible representation of the group $\operatorname{Aut}(\Delta)$ defined up to equivalence}\Bigr{)}.$
Recall that any finite group can be realized as a group of automorphisms of some finite graph.
$3^{\lozenge}.$ Consider a linear space $V$ over a finite field $\mathbb{F}$ consisting of sequences $v=(v_{1},v_{2},\dots)$, where $v_{j}\in\mathbb{F}$ and $v_{k}=0$ for sufficiently large $k$. The group $\operatorname{GL}(\infty,\mathbb{F})$ of all linear transformations of $V$ can be regarded as the group of all invertible infinite matrices over $\mathbb{F}$, which have finite number of elements in each column. The group $\operatorname{GL}(\infty,\mathbb{F})$ is perfect oligomorphic, perfect subsets are finite-dimensional linear subspaces in $V$.
Objects of the train are finite dimensional linear spaces over $\mathbb{F}$. A morphism $X\to Y$ is a partial linear bijection, i.e., an invertible operator from a subspace $A\subset X$ to a subspace $B\subset Y$. Groups of automorphisms are group of linear transformations $\operatorname{GL}(X)$. So irreducible representation are enumerated by a number $n=0$, $1$, $2$, $\dots$ and a complex irreducible representation of the group $\operatorname{GL}(n,\mathbb{F})$.
$4^{\lozenge}.$ ([38], [40]) Consider the group $\operatorname{GL}^{\diamond}(\infty,\mathbb{F})$ consisting of infinite invertible matrices over a finite field $\mathbb{F}$, which have finite number of elements in each row and each column (inverse matrices must have the same form). Let $V$ be as in the previous example, let $V^{\prime}$ be a copy of $V$. Let $\operatorname{GL}^{\diamond}(\infty,\mathbb{F})$ act in $V\cup V^{\prime}$ by the formula $v\oplus v^{\prime}\mapsto gv\oplus g^{t-1}v^{\prime}$, where t denotes the transposition. Notice that this action preserves the natural pairing $V\times V^{\prime}\to\mathbb{F}$ defined by $\{v,v^{\prime}\}=\sum v_{j}v^{\prime}_{j}$. The group $\operatorname{GL}^{\diamond}(\infty,\mathbb{F})$ is a perfect oligomorphic subgroup in the symmetric group $S(V\cup V^{\prime})$, any perfect subset in $V\cup V^{\prime}$ is a union $X\cup X^{\prime}$ of a subspace $X\subset V$ and a subspace $X^{\prime}\subset V^{\prime}$. Morphisms $(X,X^{\prime})\to(Y,Y^{\prime})$ are pairs of partial linear bijections $\mu:X\to X^{\prime}$, $\nu:Y\to Y^{\prime}$ such that for any $v\in\operatorname{dom}\mu$, $v^{\prime}\in\operatorname{dom}\nu$ we have $\{\mu(v),\nu(v^{\prime})\}=\{v,v^{\prime}\}$.
Consider a pair $(X,X^{\prime})$ of finite-dimensional linear spaces over $\mathbb{F}$ and a bilinear form $\lambda:X\times X^{\prime}\to\mathbb{F}$. We consider the subgroup $\operatorname{GL}(X\Join X^{\prime},\lambda)$ in $\operatorname{GL}(X)\times\operatorname{GL}(X^{\prime})$ preserving $\lambda$. Irreducible representations of $\operatorname{GL}^{\diamond}(\infty,\mathbb{F})$ (and of its train) are enumerated by the following data. We choose a pair $(X,X^{\prime})$ with $\lambda$ and an irreducible representation of the group $\operatorname{GL}(X\Join X^{\prime},\lambda)$.
$5^{\lozenge}.$ Let $\mathbb{F}$ be a finite field of characteristic $\neq 2$. Consider a non-degenerate quadratic form $\upsilon$ on $V$, all such forms are equivalent, so we can assume $\lambda(v)=\sum v_{j}^{2}$. Consider the group $\mathrm{O}(V)$ of linear transformations preserving this form. It is perfect oligomorphic on $V$, perfect subsets are finite-dimensional linear subspaces in $V$ (the proof for cubic forms from the next section is valid for quadratic forms). Objects of the train are finite-dimensional spaces $X$ equipped with quadratic forms. A morphisms $X\to Y$ is a bijective linear map from a subspace $A\subset X$ to a subspace $B\subset Y$ preserving the form.
2.8. Some remarks and questions. 1) In Examples $1^{\lozenge}$, $2^{\lozenge}$, $4^{\lozenge}$, $5^{\lozenge}$, we really have reduced double cosets.
For the group $K=\operatorname{Aut}({\mathbb{Q}},<)$ double cosets $K(A)\backslash K/K(B)$ are enumerated by the following data. We consider an ordered set $Q$ and two order preserving injective maps $\lambda:A\to Q$, $\mu:B\to Q$ such that $\lambda(A)\cup\mu(B)=Q$. Taking the product $\lambda^{*}\mu$ in the sense of partial bijections we get the corresponding element of $[\![K(A)\backslash K/K(B)]\!]$.
For the group $\operatorname{Aut}(\mathcal{R})$, picture is similar. Let $A$, $B$ be finite induced subgraphs of $\mathcal{R}$. Double cosets $K(A)\backslash K/K(B)$ are enumerated by the following data:
— a finite graph $Q$;
— embeddings $\lambda:A\to Q$, $\mu:B\to Q$ establishing isomorphisms of $A$ and $B$ with induced subgraphs in $Q$ such that $\lambda(A)$ and $\mu(B)$ covers the set of vertices of $Q$.
In both cases, I do not see a possibility to define a natural operation
$K(A)\backslash K/K(B)\,\times\,K(B)\backslash K/K(C)\,\to\,K(A)\backslash K/K(C),$
the multiplication arises after reduction.
For the group $\operatorname{GL}^{\diamond}(\infty,\mathbb{F})$ it is possible to assign a product for any two double cosets (it is a straightforward copy of a similar operation for real classical groups). This product was examined and clarified by Olshanski [38]. But in this case unitary representations do not separate double cosets, and we have a strong reduction.
2) As we mentioned in Introduction, theory of oligomorphic groups arose not in representation theory (see [8], [40], [4]). From the point of view of refined representation theory (as semisimple Lie groups, symmetric groups, Chevalley groups, and infinite-dimensional groups) their unitary representations can seem exotic from one hand and relatively poor from the other hand.
We have seen a classifications of irreducible unitary representations of the group $\operatorname{Aut}(\mathcal{R})$, which includes classification of all irreducible representations of all finite groups. In this sense, the situation looks very complicated and very rich. Nevertheless, visible to the author problems of representation theory for the group $\operatorname{Aut}(\mathcal{R})$ are reduced using the Mackey machinery of induced representations [22], [23], [2] to problems for finite groups (which in this generality are unsolvable). Certainly, it is nice and impressive that the group $\operatorname{Aut}(\mathcal{R})$ admits a classification of unitary representations. But it is not clear how to continue this, if we want a continuation in the representation-theoretical spirit.
3) At first glance, we have a similar picture for complete infinite symmetric group $S_{\infty}$ and the groups $\operatorname{GL}(\infty,\mathbb{F})$. Their irreducible representations are enumerated by all irreducible representations of all finite symmetric groups $S_{n}$ and all finite groups $\operatorname{GL}(n,\mathbb{F})$ respectively. Natural representation-theoretical problems for $S_{\infty}$ and $\operatorname{GL}(\infty,\mathbb{F})$ are reduced to heavy representation-theoretical problems for finite symmetric groups $S_{n}$ and finite groups $\operatorname{GL}(n,\mathbb{F})$. It seems that we do not meet new entities passing from $S_{n}$ to $S_{\infty}$ and from $\operatorname{GL}(n,\mathbb{F})$ to $\operatorname{GL}(\infty,\mathbb{F})$.
However, Lieberman classification theorem (see [20], [35], [25]) for representations of $S_{\infty}$ is an important element of a wider ’representation theory of infinite symmetric groups’ (see, for example, [35], [19], [27], [7]). This theory was initiated by the work of Elmar Thoma [39] (1964) on ’characters of infinite symmetric group’. Olshanski [35] explained that the Thoma characters correspond to unitary representations of the following ’bisymmetric group’: we consider the subgroup in $S_{\infty}\times S_{\infty}$ consisting of pairs $(g_{1},g_{2})$ such that $g_{1}g_{2}^{-1}$ are finitely supported permutations [^10] Consider an irreducible unitary representation $\rho$ of a compact group $K$ in a Euclidean space $V$. Consider the space $\mathop{\mathrm{Op}}\nolimits(V)$ of all linear operators in $V$ equipped with the inner product $\langle A,B\rangle:=\mathop{\mathrm{tr}}\nolimits AB^{*}$. The group $K\times K$ acts in $\mathop{\mathrm{Op}}\nolimits(V)$ by operators $R(k_{1},k_{2})\,A=\rho(k_{1})A\rho(k_{2})^{-1}$. The identity operator $1_{V}$ is a unique vector fixed by the diagonal subgroup $\mathop{\mathrm{diag}}\nolimits(K)\subset K\times K$. The matrix element of $1_{V}$ is $\langle R(k_{1}k_{2})1_{V},1_{V}\rangle=\mathop{\mathrm{tr}}\nolimits\rho(k_{1}k_{2}^{-1})$. It is a value of the character of $\rho$ at $(k_{1}k_{2}^{-1})$. So (in the standard terminology) ’characters of a group $K$ are functions on $K\times K$ spherical with respect to $\mathop{\mathrm{diag}}\nolimits(K)$. It turns out that a pass to infinite limit in theories of characters (and, more generally, theories of spherical functions) is more interesting than attempts of straightforward generalizations of representations (’bisymmetric group’ arises in such limit). Also, it turns out that theory of Thoma characters can be included to a wider picture that (at least now) has no counterpart for finite symmetric groups, see [27].. This group contains the diagonal consisting of pairs $(g,g)$, where $g\in S_{\infty}$. There is a big zoo of other interesting groups containing $S_{\infty}$, see [27].
Oligomorphic groups $\operatorname{GL}(\infty)$ over finite fields also can be included to a wider picture (see, e.g., [29]).
So, there is a hope that some other oligomorphic groups can be included to wider theories in representation-theoretical spirit.
4) Unitary representations of inverse limits $\lim\limits_{\leftarrow}G_{n}$ of oligomorphic groups are easily reduced to representation of prelimit oligomorphic groups $G_{n}$, see [40]. Precisely, any irreducible representation skips through one of groups $G_{n}$ (moreover, this is a general statement for any inverse limits of groups). But here we get new possibilities.
For example, consider the Bruhat–Tits tree $\mathcal{T}_{\infty}$ of infinite order, i.e., the simplicial tree whose vertices have countable valences. Fix a vertex $\xi$ and consider a ’circle’ $(\mathcal{T}_{\infty}^{n},\xi)\subset\mathcal{T}_{\infty}$ of radius $n$ with center at $\xi$. Consider the group $\operatorname{Aut}(\mathcal{T}_{\infty})$ of all automorphisms of $\mathcal{T}_{\infty}$, it subgroup $\operatorname{Aut}(\mathcal{T}_{\infty},\xi)$ fixing $\xi$, and the groups $\operatorname{Aut}(\mathcal{T}_{\infty}^{n},\xi)$. Groups $\operatorname{Aut}(\mathcal{T}_{\infty}^{n},\xi)$ are oligomorphic, the group $\operatorname{Aut}(\mathcal{T}_{\infty},\xi)$ is their inverse limit,
$\operatorname{Aut}(\mathcal{T}_{\infty},\xi)=\lim_{\leftarrow}\operatorname{Aut}(\mathcal{T}_{\infty}^{n},\xi).$
Tsankov theorem 2.1 immediately gives the classification of unitary representations of groups $\operatorname{Aut}(\mathcal{T}_{\infty}^{n},\xi)$.
An object of the corresponding category is a rooted tree $(L,\xi)$ such that distances between any vertex of $L$ and the root $\xi$ are $\leqslant n$. A morphism $(L,\xi)\to(M,\xi)$ is an isomorphism of a rooted subtree $(A,\xi)\subset(L,\xi)$ to a rooted subtree $(B,\xi)\subset(M,\xi)$.
We get a more rich situation, when we pass to natural overgroups of $\operatorname{Aut}(\mathcal{T}_{\infty},\xi)$, namely, to $\operatorname{Aut}(\mathcal{T}_{\infty})$, see [^11] The train of $\operatorname{Aut}(\mathcal{T}_{\infty})$ is not the category of finite trees and their partial isomorphisms as one might think looking to the previous examples, see [25], Sect VIII.6. [34], and to the group of spheromorphisms of $\mathcal{T}_{\infty}$, [28].
Another examples are groups $\operatorname{GL}(\infty,{\mathbb{Z}}_{p})$ and $\operatorname{GL}^{\diamond}(\infty,{\mathbb{Z}}_{p})$ over the ring ${\mathbb{Z}}_{p}$ of $p$-adic integers, they are inverse limits
(2.4) $\operatorname{GL}(\infty,{\mathbb{Z}}_{p})=\lim_{\leftarrow}\operatorname{GL}(\infty,{\mathbb{Z}}/p^{n}{\mathbb{Z}}),\quad\operatorname{GL}^{\diamond}(\infty,{\mathbb{Z}}_{p})=\lim_{\leftarrow}\operatorname{GL}^{\diamond}(\infty,{\mathbb{Z}}/p^{n}{\mathbb{Z}}).$
of oligomorphic groups. For the group $\operatorname{GL}(\infty,{\mathbb{Z}}/p^{n}{\mathbb{Z}})$ objects of the train are finite modules $L$ over the ring ${\mathbb{Z}}/p^{n}{\mathbb{Z}}$, a morphism $L\to M$ is an isomorphism of a submodule $A\subset L$ to a submodule $B\subset M$. For the groups $\operatorname{GL}^{\diamond}(\infty,{\mathbb{Z}}/p^{n}{\mathbb{Z}})$ explicit description of the trains is unknown, see [30].
Clearly, for groups (2.4) there are natural overgroups.
3. The group of automorphisms of the universal cubic spaces
3.1. The universal qubic space. Let $\mathbb{F}$ be a finite field with $q$ elements. We use the term cubic space for a linear space $V$ over $\mathbb{F}$ equipped with a symmetric trilinear (’cubic’) form $\varpi$,
$\varpi:V\times V\times V\to\mathbb{F}.$
We say that an invertible operator $V\to V$ is isocubic if it preserves the cubic form. Denote $\mathop{\mathrm{Isoc}}\nolimits(V)$ or $\mathop{\mathrm{Isoc}}\nolimits(V,\varpi)$ the group of all such maps.
Let $V$, $W$ be qubic spaces equipped with forms $\varpi_{\vphantom{|^{|}}V}$, $\varpi_{\vphantom{|^{|}}W}$. We say that a linear injective map $\iota:V\to W$ is an isocubic embedding if for any $v_{1}$, $v_{2}$, $v_{3}\in V$ we have
$\varpi_{\vphantom{|^{|}}W}\bigl{(}\iota(v_{1}),\iota(v_{2}),\iota(v_{3})\bigr{)}=\varpi_{\vphantom{|^{|}}V}(v_{1},v_{2},v_{3}).$
By $\operatorname{St}(V,W)$ (a Stiefel variety) we denote the set of all isocubic embeddings $V\to W$. The group $\mathop{\mathrm{Isoc}}\nolimits(V)\times\mathop{\mathrm{Isoc}}\nolimits(W)$ acts on $\operatorname{St}(V,W)$ in a natural way: $\iota\mapsto\sigma\iota\lambda^{-1}$, $\sigma\in\mathop{\mathrm{Isoc}}\nolimits(W)$, $\lambda\in\mathop{\mathrm{Isoc}}\nolimits(V)$.
We say that a cubic space $\mathcal{Z}$ of countable dimension satisfies the Urysohn extension property if:
for any finite-dimensional cubic space $W$ and its subspace $V$ of codimension 1, any isocubic embedding $V\to\mathcal{Z}$ extends to an isocubic embedding $W\to\mathcal{Z}$.
Theorem 3.1.
a) There exists a unique up to linear equivalence cubic space $\mathcal{Z}$ of countable dimension satisfying the Urysohn extension property.
b) The space $\mathcal{Z}$ satisfies properties:
Universality: any finite-dimensional cubic space admits an isocubic embedding in $\mathcal{Z}$;
Ultrahomogenity: For any finite-dimensional cubic space $V$ the Stiefel variety $\operatorname{St}(V,\mathcal{Z})$ is $\mathop{\mathrm{Isoc}}\nolimits(\mathcal{Z})$-homogeneous.
In particular, $\mathop{\mathrm{Isoc}}\nolimits(\mathcal{Z})$ is oligomorphic, this follows from ultrahomogeneity.
3.2. Remarks. 1) Theorem 3.1 is a very special case of Harman, Snowden [13], on cubic spaces see also [17], [5], [6]. In our degree of generality, the statement admits a trivial probabilistic proof, we need a probabilistic construction and present it in the next subsection.
2) Groups preserving symmetric or skew-symmetric bilinear forms are subjects of substantial theories. Spaces with trilinear forms are a kind of exotics. However, exceptional Lie (and Chevalley) groups admit realizations as groups preserving (symmetric or skew-symmetric) trilinear forms, see, e.g., [12], [1], [43].
3) Infinite-dimensional manifolds of embeddings of a compact manifolds $M$ in a manifold $N$ can be equipped with natural closed $k$-forms admitting infinite-dimensional groups of symmetries with infinite-dimensional stabilizers of points, see Vizman [42].
3.3. Existence of the universal cubic space. Consider a linear space $Z$ over $\mathbb{F}$ of countable dimension, fix a basis $e_{1}$, $e_{2}$, …. For each non-ordered triple of vectors we assign a random value $\zeta(e_{i},e_{j},e_{k})\in\mathbb{F}$ uniformly distributed in $\mathbb{F}$. We assume that all these values are independent. We get a random cubic form $\zeta$ on $V$. We also can say that we consider a countable direct product of copies of the field $\mathbb{F}$ equipped with a uniform measure, factors are enumerated by non-ordered triples of basis elements $(e_{i},e_{j},e_{k})$, which can coincide.
Equivalently, we take an arbitrary exhaustive sequence $X_{1}\subset X_{2}\subset\dots$ of finite dimensional subspaces such that each $X_{j}$ is equipped with a uniformly distributed cubic form. A restriction of forms from $X_{j}$ to $X_{j-1}$ sends uniform distributions to uniform distributions; we take the inverse limit.
Consider another exhaustive sequence $Y_{1}\subset Y_{2}\subset\dots$. Then each $Y_{i}$ is contained in some $X_{\alpha(i)}$ and each $X_{j}$ is contained in some $Y_{\beta(j)}$. So it produces the same distribution of cubic forms on $Z$.
In particular, the distribution of forms does not depend on a choice of a basis.
We claim that with probability 1 a form $\zeta$ satisfies the Urysohn extension property. Indeed, consider a finite dimensional cubic space $W$ and its subspace $V$ of codimension 1. Fix an isocubic embedding $\iota:V\to(Z,\zeta)$. Fix a basis $e_{1}$, …, $e_{k}$ in $\iota(V)$ and complete it to a basis $\{e_{\alpha}\}$ in $Z$. For each $\alpha>k$ we consider the subspace $W_{\alpha}$ generated by $\iota(V)$ and $e_{\alpha}$. The cubic form on $W_{\alpha}$ is determined by a collection of $k(k+1)/2+k+1$ numbers
(3.1) $\zeta(e_{i},e_{j},e_{\alpha}),\quad\zeta(e_{i},e_{\alpha},e_{\alpha}),\quad\zeta(e_{\alpha},e_{\alpha},e_{\alpha})\in\mathbb{F},$
where $i$, $j\leqslant k$. With probability $q^{-(k+1)(k+2)/2}$ this space is isocubic to $W$. Since for different $\alpha$ collections (3.1) are independent, with probability $1$ we meet a subspace isocubic to $W$.
The set of possible pairs of subspaces $V\subset W$ is countable, the set of embeddings $V\to Z$ is countable, so with probability 1 the extension of an embedding $V\to\mathcal{Z}$ is possible for any pair $V\subset W$. The Urysohn extension property is proved.
Now the universality is obvious – we take a basis $f_{1}$, …, $f_{k}$ in a cubic space $Y$, denote by $Y_{j}\subset Y$ the linear span of $f_{1}$, …, $f_{j}$. By the Urysohn property we can extend an isocubic embedding $Y_{j}\to\mathcal{Z}$ to an isocubic embedding $Y_{j+1}\to\mathcal{Z}$. We start with $Y_{0}=0$ and come to $Y_{k}=Y$.
Let us verify the ultrahomogeneity. Let $P_{k}$, $Q_{k}$ be two $k$-dimensional isocubic subspaces in a space $\mathcal{Z}$ satisfying the Urysohn property. Consider a sequence $a_{j}$ containing once all elements of $\mathcal{Z}\setminus P_{k}$ and a sequence $b_{j}$ containing once all elements of $\mathcal{Z}\setminus Q_{k}$.
Denote $P_{k+1}:=P_{k}\dotplus\mathbb{F}a_{1}$. We extend the map $\iota_{k}:P_{k}\to Q_{k}$ to an isocubic embedding $\iota_{k+1}:P_{k+1}\to\mathcal{Z}$. Denote the image by $Q_{k+1}$. Remove from the sequence $a_{j}$ all elements that are contained in $P_{k+1}$ and shift numeration of remaining elements preserving their order. Perform the same procedure with $Q_{k+1}$ and the sequence $b_{k}$.
Now we denote $Q_{k+2}=Q_{k+1}\dotplus\mathbb{F}b_{1}$ (with new $b_{1}$), take an isocubic embedding $\iota_{k+2}^{-1}:Q_{k+2}\to\mathcal{Z}$, and denote the image by $P_{k+2}$. We repeat the same operation a countable number of times, interlacing $P_{m}$ and $Q_{m}$, and come to an isocubic bijection $\mathcal{Z}\to\mathcal{Z}$ sending $P_{k}\to Q_{k}$.
It remains to verify the uniqueness of $\mathcal{Z}$. We take two cubic spaces $\mathcal{Z}_{1}$, $\mathcal{Z}_{2}$ satisfying the Urysohn property and repeat the arguments from the proof of the ultrahomogeneity starting from $P_{0}=0$ in $\mathcal{Z}_{1}$ and $Q_{0}=0$ in $\mathcal{Z}_{2}$. $\square$
Remark. This proof is a copy of the probabilistic proofs of existence of the Rado graph, see Erdős, P., Rényi [10] (see, also, [9]). The proof of universality is a copy of Urysohn’s considerations. $\boxtimes$
3.4. Some corollaries from the probabilistic construction. Let $L$ be a linear space over $\mathbb{F}$ of countable dimension with a fixed basis $e_{l}$ equipped with a cubic form $\lambda$. Let $M$ be another linear space over $\mathbb{F}$ of countable dimension with a fixed basis $f_{m}$. Denote be $\mathfrak{Z}(L)$ the space of all cubic forms $\zeta$ on [^12] We denote by $S\dotplus T$ a direct sum of linear spaces, they are not ’orthogonal’ in any sense.
$Z:=L\dotplus M$
extending the form $\lambda$ on $L$. A point $(Z,\zeta)$ of $\mathfrak{Z}(L)$ is uniquely determined by a collection of values
$\zeta(e_{l},e_{l^{\prime}},f_{k}),\quad\zeta(e_{l},f_{k},f_{k^{\prime}}),\quad\zeta(f_{k},f_{k^{\prime}},f_{k^{\prime\prime}}),$
where $l$, $l^{\prime}$, $k$, $k^{\prime}$, $k^{\prime\prime}\in{\mathbb{N}}$. We define a probabilistic measure on $\mathfrak{Z}(L)$ assuming that all these random values are uniformly distributed and independent.
Proposition 3.2.
Almost all spaces $(Z,\zeta)\in\mathfrak{Z}(L)$ are isocubic to $\mathcal{Z}$.
Proof. It is sufficient to verify the Urysohn extension property. Let $V\subset W$ be finite dimensional cubic spaces, $V$ has codimension 1 in W. Consider an embedding $\varphi:V\to L\dotplus M$, such that for any $v_{1}$, $v_{2}$, $v_{3}\in V$ satisfying $\varphi(v_{i})\in L$ we have
$\lambda\bigl{(}\varphi(v_{1}),\varphi(v_{2}),\varphi(v_{3})\bigr{)}=\varpi_{V}(v_{1},v_{2},v_{3}).$
Consider the subset $\mathfrak{Z}_{\varphi}\subset\mathfrak{Z}$ consisting of all spaces $(Z,\zeta)$ such that
$\zeta\Bigr{|}_{L}=\lambda,\qquad\zeta\bigl{(}\varphi(v_{1}),\varphi(v_{2}),\varphi(v_{3})\bigr{)}=\varpi_{V}(v_{1},v_{2},v_{3}).$
Clearly, $\mathfrak{Z}_{\varphi}$ is a subset of nonzero measure in $\mathfrak{Z}(L)$. So, almost all spaces $(Z,\zeta)\in\mathfrak{Z}_{\varphi}$ are isocubic to $\mathcal{Z}$.
Consider a basis $h_{\alpha}\in V$, let $h_{\mathrm{next}}\in W\setminus V$. Decompose vectors $\varphi(h_{\alpha})$ in the basis,
$\varphi(h_{\alpha})=\sum a_{\alpha,i}e_{i}+\sum b_{\alpha,j}f_{j}.$
Only finite number of $f_{j}$ actually enters to such sums, so all such $j$ are less than some $N$. We take a sequence of subspaces $\varphi(V)+\mathbb{F}f_{N}$, $\varphi(V)+\mathbb{F}f_{N+1}$, $\dots$ All random values
$\zeta\bigl{(}\varphi(h_{\alpha}),\varphi(h_{\beta}),f_{N+k}\bigr{)},\quad\zeta\bigl{(}\varphi(h_{\alpha}),f_{N+k},f_{N+k}\bigr{)},\quad\zeta\bigl{(}f_{N+k},f_{N+k},f_{N+k}\bigr{)}$
are independent. Therefore, a.s. there exists $k$ such that the linear map $\widetilde{\varphi}:W\to(Z,\zeta)$ defined by $\widetilde{\varphi}(h_{\alpha})=\varphi(h_{\alpha})$, $\widetilde{\varphi}(h_{\mathrm{next}})=f_{N+k}$ is isocubic. $\square$
We need the following modification of this proposition (extracted from its proof).
Let $A\subset L$ be a subspace consisting of linear combinations of some basis vectors $e_{\alpha_{1}}$, …, $e_{\alpha_{\mu}}$. Let $B\subset M$ be a subspace consisting of linear combinations of some basis vectors $f_{\beta_{1}}$, …, $f_{\beta_{\nu}}$. Fix a cubic form $\varpi_{\vphantom{\bigl{|}}A\dotplus B}$ on $A\dotplus B$ such that $\varpi_{\vphantom{\bigl{|}}A\dotplus B}\Bigr{|}_{A}=\lambda\Bigr{|}_{A}$. Consider the space $\mathfrak{Z}\bigl{(}L|A,B\bigr{)}$ of all cubic forms $\zeta$ on $L\dotplus M$ such that
$\zeta\Bigr{|}_{A\dotplus B}=\varpi_{\vphantom{\bigl{|}}A\dotplus B},\qquad\zeta\Bigr{|}_{L}=\lambda.$
Such form is uniquely determined by collection of values
(3.2) $\zeta(e_{i},e_{i^{\prime}},f_{j}),\quad\zeta(e_{i},f_{j},f_{j^{\prime}}),\quad\zeta(f_{j},f_{j^{\prime}},f_{j^{\prime\prime}}),$
where at least one $i$, $i^{\prime}\notin\{\alpha_{1},\dots,\alpha_{\mu}\}$ or at least one $j$, $j^{\prime}$, $j^{\prime\prime}\notin\{\beta_{1},\dots,\beta_{\nu}\}$. We define the probabilistic measure on $\mathfrak{Z}\bigl{(}L|A,B\bigr{)}$ assuming that all random values (3.2) are independent and uniformly distributed.
Corollary 3.3.
Almost all spaces $(Z,\zeta)\in\mathfrak{Z}\bigl{(}L|A,B\bigr{)}$ are isocubic to $\mathcal{Z}$.
Proof. The space $\mathfrak{Z}\bigl{(}L|A,B\bigr{)}$ is a subset of nonzero measure in $\mathfrak{Z}(L)$. $\square$
3.5. Open subgroups of the group $\mathop{\mathrm{Isoc}}\nolimits(\mathcal{Z})$. For any finite-dimensional subspace $V\subset\mathcal{Z}$ we denote by $K(V)\subset\mathop{\mathrm{Isoc}}\nolimits(\mathcal{Z})$ the subgroup fixing all elements of $V$, by $K^{\circ}(V)\subset\mathop{\mathrm{Isoc}}\nolimits(\mathcal{Z})$ the subgroup sending $V$ to itself. Clearly,
$K^{\circ}(V)/K(V)\simeq\mathop{\mathrm{Isoc}}\nolimits(V).$
The group $\mathop{\mathrm{Isoc}}\nolimits(\mathcal{Z})$ is a closed subgroup in the symmetric group $S(\mathcal{Z})$, the subgroups $K(V)\subset\mathop{\mathrm{Isoc}}\nolimits(\mathcal{Z})$ are open and form a fundamental system of neighborhoods of the unit.
Proposition 3.4.
All finite-dimensional subspaces in $\mathcal{Z}$ are perfect subsets.
Theorem 3.5.
For any open subgroup $H\subset\mathop{\mathrm{Isoc}}\nolimits(\mathcal{Z})$ there is a finite-dimensional subspace $V$ such that $K(V)\subset H\subset K^{\circ}(V)$.
Thus, the subgroup $\mathop{\mathrm{Isoc}}\nolimits(\mathcal{Z})\subset S(\mathcal{Z})$ is perfect oligomorphic, perfect subsets are linear subspaces in $\mathcal{Z}$. So, Theorem 2.1 gives us a classification of unitary representations. By Theorem 2.5, we have a well-defined category of reduced double cosets, it is equivalent to the category of partial linear isocubic bijections.
Corollary 3.6.
Objects of the category $\operatorname{Tra}\bigl{(}\mathop{\mathrm{Isoc}}\nolimits(\mathcal{Z})\bigr{)}$ are finite-dimensional cubic linear spaces. A morphism $V\to W$ is an isocubic linear bijection of a linear subspace $A\subset V$ to a linear subspace $B\subset W$.
3.6. Proof of Proposition 3.4.
Lemma 3.7.
Let $Y$, $V\subset\mathcal{Z}$ be finite-dimensional subspaces, $V\cap Y=0$. Then there exists a countable collection of $g_{j}\in K(V)$ such that the sum
$L:=V+Y+g_{1}Y+g_{2}Y+\dots$
is direct and has countable codimension in $\mathcal{Z}$
Proposition 3.4 follows from the lemma. It is sufficient to consider one-dimensional $Y$.
Proof of Lemma 3.7. We apply Proposition 3.2. Let $Y^{1}$, $Y^{2}$, $\dots$ be disjoint copies of $Y$. Consider a direct sum
(3.3) $L:=V\dotplus Y\dotplus Y^{1}\dotplus Y^{2}\dotplus\dots.$
We define cubic form on each subspace $V\dotplus Y\subset L$, $V\dotplus Y^{j}\subset L$ assuming that tautological maps from the initial space $V\dotplus Y$ to the corresponding subspace $V\dotplus Y\subset L$ and the subspaces $V\dotplus Y_{j}\subset L$ are isocubic. We consider an arbitrary cubic form on $L$ compatible with the forms on $V\dotplus Y$, $V\dotplus Y^{j}$. Next, we consider the corresponding space $\mathfrak{Z}(L)$ and take an element $(Z,\zeta)\in\mathfrak{Z}(L)$ isomorphic to $\mathcal{Z}$. By the ultrahomogeneity, we have $g_{j}\in\mathop{\mathrm{Isoc}}\nolimits(Z,\zeta)$ satisfying desired conditions. $\square$
3.7. Proof of Theorem 3.5. Let $H$ be an open subgroup, $H\supset K(V)$, and there exists $\xi\in H\setminus K^{\circ}(V)$. It is sufficient to prove that there exists a proper subspace $W\subset V$ such that $K(W)\subset H$.
Notice that $\xi K(V)\xi^{-1}=K(\xi V)$ and $\xi V\neq V$. So it is sufficient to prove the following statement (assuming $V^{\prime}:=\xi V$):
Lemma 3.8.
Let a subgroup $H\subset\mathop{\mathrm{Isoc}}\nolimits(\mathcal{Z})$ contain $K(V)$ and $K(V^{\prime})$. Then it contains $K(V\cap V^{\prime})$.
Lemma 3.9.
In the notation of Lemma 3.8, denote $W:=V\cap V^{\prime}$. Decompose $V=W\dotplus X$, $V^{\prime}=W\dotplus Y$. Let $W\dotplus\widetilde{X}\dotplus\widetilde{Y}$ be a subspace in $\mathcal{Z}$ such that there is an isocubic map $r:W\oplus X\oplus Y\to W\oplus\widetilde{X}\oplus\widetilde{Y}$ identical on $W$ and sending $X\to\widetilde{X}$, $Y\to\widetilde{Y}$. Then there is an element $h\in H$ such that the restriction of $h$ to $W\oplus X\oplus Y$ is $r$.
Proof of Lemma 3.9. We have $K(W\oplus X)$, $K(W\oplus Y)\subset H$. Notice that
$h\in H,\quad K(A)\subset H\qquad\Rightarrow\qquad K(hA)=hK(A)h^{-1}\subset H.$
Step 1. Let us show that without loss of generality we can assume that
(3.4) $\widetilde{X}\cap\bigl{(}W\dotplus X\dotplus Y\bigr{)}=0.$
By Lemma 3.7 and ulrahomogeneity, there is a countable collection of maps $g_{j}\in K(W\dotplus Y)$ such that a sum
$W+Y+X+g_{1}X+g_{2}X+\dots$
is direct. For sufficiently large $N$, we have $\widetilde{X}\cap(W\dotplus g_{N}X)=0$. Now we can replace $X\mapsto g_{N}X$, $r\mapsto rg_{N}^{-1}$. So, we can assume
$\widetilde{X}\cap(W\dotplus X)=0.$
We again apply Lemma 3.7. There exists a countable collection of maps $h_{j}\in K(W\dotplus X)$ such that a sum $W+X+Y+h_{1}Y+h_{2}Y+\dots$ is direct. For sufficiently large $N$, we have $\widetilde{X}\cap(W\dotplus X\dotplus h_{N}Y)=0$. Now we can replace $Y\mapsto h_{N}Y$, $r\mapsto rh_{N}^{-1}$. We come to the case (3.4).
Step 2. We once again apply Lemma 3.7 and keep in mind its proof. Choose bases
(3.5) $w_{\alpha}\in W,\quad x_{\beta}\in X,\quad y_{\gamma}\in Y.$
Consider the space $L$ as in the proof of Lemma 3.7 assuming $V=W\dotplus X$. Consider copies $y_{\gamma}^{j}\in Y^{j}$ of vectors $y_{\gamma}$. Such vectors together with (3.5) form a basis in the space $L$. So we get a space $L$ equipped with a cubic form, say $\lambda$.
Next, consider the space $Z:=L\dotplus M$ as in Proposition 3.2. Let us identify the basis $rx_{\alpha}\in\widetilde{X}$ with first basis elements $f_{1}$, …, $f_{s}\in M$. Consider the restriction of our initial cubic form $\zeta$ to $W\dotplus X\dotplus Y\dotplus\widetilde{X}$, and assign this cubic form to the corresponding subspace in $L\dotplus M$.
We came to an object described in Corollary 3.3, we have fixed cubic forms on $L$ and $W\dotplus X\dotplus Y\dotplus\widetilde{X}$ and these forms are compatible on their intersection $W\dotplus X\dotplus Y$. So, we get a probabilistic space $\mathfrak{Z}(L|W\dotplus X\dotplus Y,\widetilde{X})$ whose points are forms $\zeta$ on $Z:=L\dotplus M$ extending our forms to the whole space. By Corollary 3.3, a cubic space $(Z,\zeta)$ a.s. is isomorphic to $\mathcal{Z}$. In particular, this space contains a copy of $\widetilde{Y}$ such that a new $W\dotplus X\dotplus Y\dotplus\widetilde{X}+\widetilde{Y}$ is isocubic to the initial $W\dotplus X\dotplus Y\dotplus\widetilde{X}+\widetilde{Y}$. Position of new $\widetilde{Y}$ depends on a random form $\zeta$, choice of this $\widetilde{Y}$ is not unique (it can be done in a canonical way, but we do not need this).
Let $j$ range in $1$, $2$, $3$, …, and $w_{\alpha}$, $x_{\beta}$, $y_{\gamma}$ range in sets of basis vectors in $W$, $X$, $Y$. Then the following collection of $\mathbb{F}$-valued random variables
$\zeta(y_{\gamma}^{j},y_{\gamma^{\prime}}^{j},rx_{\beta}),\quad\zeta(y_{\gamma}^{j},rx_{\beta},rx_{\beta^{\prime}}),\quad\zeta(y_{\gamma}^{j},w_{\beta},rx_{\beta})$
is independent. Therefore, with probability 1 there is $j_{*}$ such that the following (finite) collection of identities holds:
(3.6) $\zeta(y_{\gamma}^{j_{*}},y_{\gamma^{\prime}}^{j_{*}},rx_{\beta})$ $=\zeta(y_{\gamma}^{j_{*}},y_{\gamma^{\prime}}^{j_{*}},x_{\beta});$
(3.7) $\zeta(y_{\gamma}^{j_{*}},rx_{\beta},rx_{\beta^{\prime}})$ $=\zeta(y_{\gamma}^{j_{*}},x_{\beta},x_{\beta^{\prime}});$
(3.8) $\zeta(y_{\gamma}^{j_{*}},w_{\beta},rx_{\beta})$ $=\zeta(y_{\gamma}^{j_{*}},w_{\beta},x_{\beta}).$
Consider the map
$\sigma:W\oplus X\oplus Y^{j_{*}}\to W\oplus rX\oplus Y^{j_{*}},$
which is identical on $W\oplus Y^{j_{*}}$ and equals $r$ on $X$ (and therefore $\sigma$ equals $r$ on $W\oplus X$). This map is $\zeta$-isocubic. Indeed,
1) by definition, it is isocubic on $W\oplus Y^{j_{*}}$;
2) $r$ is isocubic on $W\oplus X$;
3) by (3.6)-(3.7) it is isocubic on $X\oplus Y^{j_{*}}$;
4) by (3.8) it preserves the trilinear form on $Y^{j_{*}}\times X\times W$.
The map $\sigma$ extends to an element $\widetilde{\sigma}$ of the subgroup $K(W\oplus Y^{j_{*}})\subset H$. Conjugating $K(W\oplus X)$ by $\widetilde{\sigma}$, we get that $K(W\oplus\widetilde{X})\subset H$. Also, by the construction, we have the following isocubic maps
$W\dotplus\widetilde{X}\dotplus\widetilde{Y}\,\leftarrow\,W\dotplus X\dotplus Y\,\rightarrow\,W\dotplus X\dotplus Y_{j^{*}}\,\rightarrow\,W\dotplus\widetilde{X}\dotplus Y_{j^{*}}.$
So, we have an isocubic map $W\oplus\widetilde{X}\oplus Y^{j_{*}}\to W\oplus\widetilde{X}\oplus\widetilde{Y}$. By the construction, it fixes $W\oplus\widetilde{X}$. Therefore there is an element $\xi$ of $K(W\oplus\widetilde{X})$ sending one subspace to another. We take $h=\xi\widetilde{\sigma}$square
Proof of Lemma 3.8. Let $r\in H$. By Lemma 3.9 there exists $h\in H$ such that $h=r$ on $V+V^{\prime}=W\dotplus X\dotplus Y$. So $h^{-1}r\in K(V+V^{\prime})$. But $K(V+V^{\prime})\subset K(V)$. $\square$
3.8. Compactification of $\mathop{\mathrm{Isoc}}\nolimits(\mathcal{Z})$. Consider the following semigroup $\Gamma(\mathcal{Z})$. Its elements are partial linear isocubic bijections from a subspace $A\subset\mathcal{Z}$ to a subspace $B\subset\mathcal{Z}$. The product is the product of partial bijections. A sequence $\gamma_{j}\in\Gamma(\mathcal{Z})$ converges to $\gamma\in\Gamma(\mathcal{Z})$ if for each $v$, $w\in\mathcal{Z}$ the condition ’$\gamma v=w$’ is equivalent to the condition ’$\gamma_{j}v=w$ for sufficiently large $j$’.
Corollary 3.10.
a) The $\Gamma(\mathcal{Z})$ is a compact semigroup with separately continuous product.
b) The group $\mathop{\mathrm{Isoc}}\nolimits(\mathcal{Z})$ is dense in $\Gamma(\mathcal{Z})$.
c) Each unitary representation of $\mathop{\mathrm{Isoc}}\nolimits(\mathcal{Z})$ extends to a weakly continuous continuous representation of $\Gamma(\mathcal{Z})$.
In particular, for each unitary representation $\rho$ of the group $\mathop{\mathrm{Isoc}}\nolimits(\mathcal{Z})$, the closure $\overline{\rho\bigl{(}\mathop{\mathrm{Isoc}}\nolimits(\mathcal{Z})\bigr{)}}$ of $\rho\bigl{(}\mathop{\mathrm{Isoc}}\nolimits(\mathcal{Z})\bigr{)}$ in the weak operator topology coincides with the image of $\Gamma(\mathcal{Z})$.
Proof. a) The semigroup of all partial bijections of a countable set is a compact semigroup with separately continuous multiplication, see, e.g., [25], Sect. VIII.1. It is easy to see that $\Gamma(\mathcal{Z})$ is closed in the semigroup of all partial bijection.
b) Consider a partial isocubic bijection $\theta\in\Gamma(\mathcal{Z})$, let $V:=\operatorname{dom}\theta$. Consider a sequence of finite-dimensional subspaces $V_{1}\subset V_{2}\subset V_{3}\subset\dots$ such that $V=\cup V_{j}$. Consider a sequence $x_{1}$, $x_{2}$, …consisting of all elements of the set $\mathcal{Z}\setminus V$. We choose a sequence $g_{k}\in\mathop{\mathrm{Isoc}}\nolimits(\mathcal{Z})$ such that for each $k$:
$g_{k}\Bigr{|}_{V_{k}}=\theta\Bigr{|}_{V_{k}}$;
— for each $j<k$, $i<k$ we have $g_{k}x_{i}\neq g_{j}x_{i}$.
This is possible due to Neumann Lemma 2.4. So, for each $x_{m}\in\mathcal{Z}\setminus V$ the sequence $g_{j}x_{m}$ consists of pairwise different elements. On the other hand, each $v\in V$ is contained in $V_{k}$ with sufficiently large $k$, so $g_{k}v=\theta v$. Therefore, $g_{k}$ converges to $\theta$.
c) It is sufficient to verify this for irreducible representations. We use the model of irreducible representations as Subsect. 2. Let $A\subset\mathcal{Z}$ be a finite-dimensional subspace, $\sigma$ an irreducible representation of $\mathop{\mathrm{Isoc}}\nolimits(A)$. The space $H(A,\sigma)$ is defined as Subsect. 2. Operators $\tau(\theta)$ in the space $H(A,\sigma)$ are defined by
$\tau(\theta)F(\iota)=\begin{cases}F(\theta^{*}\iota),\quad&\text{if}\quad A\subset\operatorname{im}\theta,\,\theta^{*}\iota A\subset\operatorname{dom}\theta;\\ 0,\quad&\text{otherwise.}\end{cases}$
3.9. Final remarks. In our considerations, symmetric trilinear forms can be replaced by skew-symmetric trilinear form, and more generally by symmetric or skew-symmetric $n$-linear forms. We also can equip a linear space with a finite collection of such forms.
Apparently, our Theorem 3.5 can be extended to arbitrary contravariant irreducible tensors over finite fields as in [13] (a non-obvious place is an extension of Proposition 3.5). Apparently, construction of [25], Sect.III.3 allows to consider also exterior forms of middle dimension.